OpenStudy (anonymous):

derivative of 9sinh(lnt). I thought it should be (9cosh(lnt))/t but apparently that is incorrect...please help!!

OpenStudy (irishboy123):

go back to the definition of sinh -- > sinh x = 1/2 (e ^ x - e^ -x) sub into that and see where you go with e^ ln t etc

OpenStudy (anonymous):

Thanks! Didn't think to do that

OpenStudy (anonymous):

I got 9/2(2t) before taking the derivative, and then just 9 after the derivative but that was wrong still

OpenStudy (zarkon):

your original answer is correct

OpenStudy (anonymous):

not according to the online hw website. It marked it as wrong

OpenStudy (zarkon):

your answer is correct...but it can be simplified. Maybe then will the website understand your answer

OpenStudy (zarkon):

\[\frac{9\cosh(\ln(t))}{t}\] \[=\frac{9}{2}\left[\frac{e^{\ln(t)}+e^{-\ln(t)}}{t}\right]\] \[=\frac{9}{2}\left[\frac{t+\frac{1}{t}}{t}\right]\] \[=\frac{9}{2}\left[1+\frac{1}{t^2}\right]\]

OpenStudy (anonymous):

I tried that too and it didn't work...

OpenStudy (zarkon):

then your website doesn't work ;)

OpenStudy (anonymous):

I know that is the answer and all that...I'm probably just going to talk to my teacher about it

OpenStudy (zarkon):

just know that what you did was correct. The website is only as good as the people that programmed it.